Claim: The standard deviation of pulse rates of adult males is more than 12 bpm. For a random sample of 129 adult males, the pulse rates have a standard deviation of 12.3 bpm. Find the value of the test statistic.
Answer #1

Here

Claim: The standard deviation of pulse rates of adult males is more than 12 bpm.

H0 : σ = 12 bpm

Ha : σ > 12bpm

sample size = n = 129

sample standard deviation = s = 12.3 bpm

here test statistic

= (n-1) s2/

= (129 -1) * 12.3 * 12.3/122 = 134.48

= 134.48

Test statistic = 134.48